The steps for finding 20 x 23 are shown. fill in the blanks
20 x23= 20x(20+___)=20x___+20x___=___+60=

Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

The steps for finding 20 x 23 are shown. fill in the blanks
20 x23= 20x(20+3)=20x20+20x3=400+60= 460
==================
Cheers,
Stan H.
================

Answer by jim_thompson5910(35256)

20 x 23= 20x(20+_3_)=20x_20_+20x_3_=_400_+60 = _460_